# MOSAIC: Generating Consistent, Privacy-Preserving Scenes from Multiple Depth Views in Multi-Room Environments

### v1

@@ 1 Introduction
Various approaches for scene level multiview image generation have been proposed, with autoregressive generation techniques [12, 7, 41, 40] representing the current mainstream trend. These approaches leverage powerful inpainting models to iteratively render unseen parts of the scene from sequential camera perspectives. However, these approaches exhibit style drift, where images generated for initial viewpoints often differ stylistically from those produced when revisiting the same location from different angles, creating visual inconsistencies. Moreover, warp-and-inpainting methods frequently encounter depth misalignment issues, leading to compounding errors over sequential generations [32, 20]. Unlike prior methods that perform diffusion synchronization tasks [2, 17, 8, 19, 15] in constrained settings (limited viewpoint variations), our task requires coherence between viewpoints based on arbitrary trajectory features, extensive perspective changes, and spatial discontinuities. This fundamental distinction makes existing methods ineffective for complex multi-room environments with widely-varying camera positions.
@@ 2 Related Work
Diffusion-Guided Scene-Level Multi-view Generation. Recent advances in 2D diffusion-based generative models [31, 10, 29, 42] have catalyzed significant progress in scene-level multiview image generation—particularly in the use of autoregressive generation techniques that leverage warping and inpainting mechanisms for, e.g., scene completion [32, 18, 30]. These mechanisms can introduce style drift errors where images in later iterations differ stylistically from the initial ones, leading to artifacts and distortions. Alternative approaches attempt to maintain style consistency by generating panoramic representations using off-the-shelf diffusion models [32, 18, 30, 36]. While these methods achieve internal coherence, they cannot handle multiple rooms with complex layouts or flexible camera trajectories, limiting their practicality in the real world. Recent video diffusion models demonstrate promising capabilities for generating temporally consistent image sequences  [4, 11, 39, 44]. However, they do not apply to our task due to the inability to incorporate depth conditions and limited tolerance against perspective changes. Our method, on the other hand, explicitly models cross-view consistency and generalizes to arbitrary viewpoints without imposing layout constraints.
@@ 4.1 Preliminaries
where $\sigma\in\mathbb{R}_{\geq 0}^{T}$ is a real vector and $z_{1:T}$ are latent variables in $\mathcal{Z}$. The forward process $q_{\sigma}\left(z_{t-1}\mid z_{t},z_{0}\right)$ is parameterized by a decreasing sequence $\alpha_{1:T}\in(0,1]^{\top}$. To approximate $q(z_{0})$, DDIM learns a generative process
@@ 4.2 MOSAIC Formulation
Intuitively, the sampling procedure in Eq. 6 would generate inconsistent RGB views because the generative process $p_{\theta}$ is trained with independent samples $z_{0}\sim q(z_{0})$. In our case, however, the samples $z_{0}^{[1:N]}$ are indeed dependent since they are taken from the same complete scene ${\bar{z}_{0}}$. To fundamentally address the consistency issue, our key insight is to explicitly model such dependency by incorporating extra projection-based conditionals $q(z_{0}^{[i]}\mid{\bar{z}_{0}})$ in the forward procedures. In practice, given ${\bar{z}_{0}}$, $z_{0}^{[i]}$ can be collected by simply sampling a camera pose with projection function $\pi^{[i]}$ and setting $z_{0}^{[i]}=\pi^{[i]}({\bar{z}_{0}})$; subsequent latents $z_{1:T}^{[i]}$ can be sampled as usual. The updated forward procedure for $N$ depth views from the same scene ${\bar{z}_{0}}$ can be rewritten from Eq. Eq. 2 as:

### v2

@@ 1 Introduction
Various approaches for scene level multiview image generation have been proposed, with autoregressive generation techniques [13, 8, 42, 41, 7] representing the current mainstream trend. These approaches leverage powerful inpainting models to iteratively render unseen parts of the scene from sequential camera perspectives. However, these approaches exhibit style drift, where images generated for initial viewpoints often differ stylistically from those produced when revisiting the same location from different angles, creating visual inconsistencies. Moreover, warp-and-inpainting methods frequently encounter depth misalignment issues, leading to compounding errors over sequential generations [33, 21]. Unlike prior methods that perform diffusion synchronization tasks [2, 18, 9, 20, 16] in constrained settings (limited viewpoint variations), our task requires coherence between viewpoints based on arbitrary trajectory features, extensive perspective changes, and spatial discontinuities. This fundamental distinction makes existing methods ineffective for complex multi-room environments with widely-varying camera positions.
@@ 2 Related Work
Diffusion-Guided Scene-Level Multi-view Generation. Recent advances in 2D diffusion-based generative models [32, 11, 30, 43] have catalyzed significant progress in scene-level multiview image generation—particularly in the use of autoregressive generation techniques that leverage warping and inpainting mechanisms for, e.g., scene completion [13, 8, 42, 41, 7]. These mechanisms can introduce style drift errors where images in later iterations differ stylistically from the initial ones, leading to artifacts and distortions. Alternative approaches attempt to maintain style consistency by generating panoramic representations using off-the-shelf diffusion models [33, 19, 31, 37]. While these methods achieve internal coherence, they cannot handle multiple rooms with complex layouts or flexible camera trajectories, limiting their practicality in the real world. Recent video diffusion models demonstrate promising capabilities for generating temporally consistent image sequences  [4, 12, 40, 45]. However, they do not apply to our task due to the inability to incorporate depth conditions and limited tolerance against perspective changes. Our method, on the other hand, explicitly models cross-view consistency and generalizes to arbitrary viewpoints without imposing layout constraints.
@@ 4.1 Preliminaries
where $\sigma\in\mathbb{R}_{\geq 0}^{T}$ is a real vector and $z_{1:T}$ are latent variables in $\mathcal{Z}$. The inference process $q_{\sigma}\left(z_{t-1}\mid z_{t},z_{0}\right)$ is parameterized by a decreasing sequence $\alpha_{1:T}\in(0,1]^{\top}$. To approximate $q(z_{0})$, DDIM learns a generative process
@@ 4.2 MOSAIC Formulation
Intuitively, the sampling procedure in Eq. 6 would generate inconsistent RGB views because the generative process $p_{\theta}$ is trained with independent samples $z_{0}\sim q(z_{0})$. In our case, however, the samples $z_{0}^{[1:N]}$ are indeed dependent since they are taken from the same complete scene ${\bar{z}_{0}}$. To fundamentally address the consistency issue, our key insight is to explicitly model such dependency by incorporating extra projection-based conditionals $q(z_{0}^{[i]}\mid{\bar{z}_{0}})$ in the inference procedures. In practice, given ${\bar{z}_{0}}$, $z_{0}^{[i]}$ can be collected by simply sampling a camera pose with projection function $\pi^{[i]}$ and setting $z_{0}^{[i]}=\pi^{[i]}({\bar{z}_{0}})$; subsequent latents $z_{1:T}^{[i]}$ can be sampled as usual. The updated inference procedure for $N$ depth views from the same scene ${\bar{z}_{0}}$ can be rewritten from Eq. 2 as: